flawdababynai:

A bicycle manufacturing company makes a particular type of bike. Each child bike requires 4 hours to build and 4 hours to test. Each adult bike requires 6 hours to build and 4 hours to test. With the number of workers, the company is able to have up to 120 hours of building time and 100 hours of testing time for a week. If c represents child bikes and a represents adult bikes, can the company build 20 child bikes and 6 adult bikes in a week.

flawdababynai:

please help

Extrinix:

Ok so, \(4~hrs~to~build~,~4~hrs~to~test\) = 8 hrs in total for a child's bike \(6~hrs~to~build~,~4~hrs~to~test\) = 10 hrs in total for an adult bike so for this you would divide the amount of hours of building (let's say \(b\)) by the amount of hours to build a childs bike (\(c\)) + an adult bike (\(a\)) \(\frac{b}{c+a}\) now, enter the information in \(\frac{120}{4+6}\) simplify \(\frac{120}{10}\) = __ do the same for the amount of testing, \(\frac{100}{4+4}\) \(\frac{100}{8}\) now figure those out, then multiply the hours to build answer by 26 and see if it ends up less than 120 \(build~time \times 26<120\) and then do the mount of testing \(test~time \times 26<100\) if both of those are lower then or equal to 120 and 100 then the bikes can be completed, if not, they cant be completed.